OpenStudy (anonymous):

write an equation in standard form for a line with a slope of -1 passing through c(2,1)

OpenStudy (anonymous):

For this you could use slope intercept form. That is y-y1=m(x-x1) where y1 and x1 represent the point c (2,1) and m=slope which is -1. Plug those numbers in and show me what you get. From there I can help you walk through how to convert it into standard form

OpenStudy (anonymous):

y-1=-1(x-2)

OpenStudy (anonymous):

y-1=2x

OpenStudy (anonymous):

@Audeezy

OpenStudy (anonymous):

that -1 in -1(x-2) is going to be distributed. W

OpenStudy (anonymous):

y-1=1x-2

OpenStudy (anonymous):

You're only distributing the 1, not the negative in front of it. That 1 doesn't need to be written as just a " - " is understood to be -1. So when you multiply x by -1 what do you get? Same as when you muliply -2 by -1?

OpenStudy (anonymous):

y-1=-1x-2?

OpenStudy (anonymous):

@Audeezy

OpenStudy (anonymous):

when you multiply 2 negatives you get a positive. so -2 x -1 will give you 2, but since its negative times a negative you will get a positive. That would give you y-1=-x+2

OpenStudy (anonymous):

The standard form of an equation is Ax + By = C So you want X and Y on the same side and C by itself, You Also want X to be positive, so combine like terms and get x and y on 1 side where X is positive and let me know what you get

OpenStudy (anonymous):

i got y-x=2

OpenStudy (anonymous):

@Audeezy